# Model-free source seeking of exponentially convergent unicycle: theoretical and robotic experimental results

### v2

@@ Abstract
This paper introduces a novel model-free, real-time unicycle-based source seeking design. This design autonomously steers the unicycle dynamic system towards the extremum point of an objective function or physical/scalar signal that is unknown expression-wise, but accessible via measurements. A key contribution of this paper is that the introduced design converges exponentially to the extremum point of objective functions (or scalar signals) that behave locally like a higher-degree power function (e.g., fourth-degree polynomial function) as opposed to locally quadratic objective functions, the usual case in literature. We provide theoretical results and design characterization, supported by a variety of simulation results that demonstrate the robustness of the proposed design, including cases with different initial conditions and measurement delays/noise. Also, for the first time in the literature, we provide experimental robotic results that demonstrate the effectiveness of the proposed design and its exponential convergence ability.
@@ 1 Introduction
Extremum seeking control (ESC) techniques are model-free, real-time control methods that drive a system to the optimum of a given objective function (ariyur2003real; review; guay2015time; scheinker2024100), which may be unknown, provided it can be measured. ESC schemes are particularly attractive across various fields due to their minimal information requirements: they rely solely on applying perturbations to system parameters or inputs, and measuring the corresponding outputs of the objective function. Using these measurements in a feedback loop, ESC algorithms iteratively adjust the system input or parameter values to drive it towards the extremum, as can be seen in many applications, e.g., krstic2008extremum; dochain2011extremum; calli2012comparison; ECC_2024; elgohary2025model; elgohary2025extremum; elgohary2025hovering; pokhrel2022novel; eisa2023; grushkovskaya2018family; grushkovskaya2024design.
@@
One conventional problem that has increasingly involved the use of ESC methods is source-seeking (bajpai2024investigating; ghods2011extremum; zhu2013cooperative; li2014cooperative; bulgur2018light). Source-seeking is the problem of steering a system (e.g., a mobile robot) autonomously to the maximum or minimum intensity of a physical/scalar signal present in a given domain using only on-board sensor measurements (e.g., heat, light, chemical concentration, among others). The advantage of ESC lies in the fact that it does not require knowledge of the signal’s distribution or explicit modeling of the system to drive the system, autonomously, to the extremum point. Due to the model-free nature of ESC methods, source-seeking techniques based on ESC utilized a simple, basic unicycle model to represent the system within the control design (pokhrel2024extremum; matveev2011navigation; khong2014multi; cochran20093; grushkovskaya2018family; suttnerkrisitic2023unicycletorque; yilmaz2025unbiased; todorovski2024newton; elgohary2025model). As a result, unicycle-based methods for source-seeking often resort to Lie bracket averaging (durr2013lie; scheinker2017bounded; grushkovskaya2018class; ghadiri2020normalized; pokhrel2023higher) in theoretical analysis and design. It is worth noting that the aforementioned unicycle-based designs for source-seeking are based on ESC laws that are gradient-based, which is more effective (e.g., guarantee exponential convergence) if the objective function or the scalar signal behaves locally like a quadratic function.
@@
On the other hand, optimization problems involving fourth-order polynomial functions have become an increasingly prevalent topic in the literature (qi2004global; zhang2015optimality). In fact, some emerging applications in science and engineering involve the need to optimize and control a fourth-order polynomial function, such as but not limited to the distribution of the coefficient of performance for power optimization ($C_{p}$) in wind turbines (eisa2019modeling) and pure-quartic solitons in optics, which exhibit fourth-order dispersion and have possible applications with lasers (soltani2023pure).
@@
Motivation and Contribution. In our recent work grushkovskaya2025extremum, new ESC laws were derived based on higher-order Lie bracket averaging. This paved the way for the prospect of having exponentially convergent ESC laws for objective functions that behave locally not as quadratic functions, but as power functions with higher-order degrees. Inspired by the prospect of realizing exponential convergence for high-order objective functions, we seek to utilize the ESC method in grushkovskaya2025extremum and propose a novel, first-of-its-kind, exponentially convergent unicycle-based design for source-seeking. In this preliminary study, we only focus on source-seeking of objective functions that behave locally as a fourth-degree power function. We prove stability of the proposed design and provide a variety of simulation results that demonstrate the robustness of the proposed design, including cases with different initial conditions and measurement delays/noise. Additionally, we provide experimental results using TurtleBot3 robot (similar to ECC_2024; elgohary2025model) and we compare our design (third-order Lie bracket-based) with traditional results using first-order Lie bracket-based design. We also perform a light source-seeking experiment in a complete model-free fashion. Our results illustrate the potential and effectiveness of the proposed exponentially convergent unicycle design for source-seeking.
@@ 2 Brief Background
In this section, we provide brief context regarding the exponentially convergent ESC law that we use for higher-order objective functions. We consider the two-input control-affine ESC of the following form, as introduced in grushkovskaya2025extremum:
@@
where $x\in D\subset\mathbb{R}^{n}$ denotes the system’s state, $J:D\to\mathbb{R}$ is the objective function, $u_{1}^{\varepsilon}(t)=\varepsilon^{1/N-1}v_{1}(t/\varepsilon)$, $u_{2}^{\varepsilon}(t)=\varepsilon^{1/N-1}v_{2}(t/\varepsilon)$, $N\in\mathbb{N}$ denotes the order of the Lie bracket to be excited, and $\varepsilon>0$ is a small positive parameter, such that the dithers $v_{1}(t/\varepsilon)$ and $v_{2}(t/\varepsilon)$ excite the Lie bracket
@@
Moreover, we adopt all mathematical assumptions on the vector fields $g_{1}$, $g_{2}$ provided in (grushkovskaya2025extremum, Section II) and recall that $g_{1},g_{2}:\mathbb{R}\to\mathbb{R}^{n}$ are smooth vector fields evaluated at the scalar argument $J(x)$. However, we state below the assumption clarifying the structure and class of objective functions admitted in grushkovskaya2025extremum as well as in this paper:
@@
Assumption A1.1 states that $J$ has an isolated minimum, $x^{*}$, at which it has a value $J^{*}$. Assumption A1.2 provides the requirement that the admitted objective function, $J$, behaves locally as a power function of degree $m$; this is consistent with ESC literature (e.g., grushkovskaya2018class).
The selection of the dither signals is done as in grushkovskaya2024design; grushkovskaya2025extremum, conditions C1–C3:
@@
From grushkovskaya2025extremum, $g_{1},\ g_{2}$ are chosen according to:
@@
with the simple selection to satisfy (2) being $g_{1}(z)=-1^{(N+1)}z$ and $g_{2}(z)=1$. It is also to be noted that the conditions C1–C3 above guarantee that only the desired Lie bracket is excited while all others vanish per Gau14; pokhrel2023higher; grushkovskaya2024design. In this paper, we are interested in third-order Lie bracket excitation (i.e., condition C3).
@@ 3.1 Proposed Design
We aim to use an ESC law based on condition C3, provided in the previous section, to achieve an exponentially convergent unicycle design. Let us consider the kinematic differential equations for unicycle dynamics with constant angular velocity as follows (elgohary2025model; grushkovskaya2018family):
@@
where $x$ and $y$ describe the current position/coordinates, $\mathrm{v}$ is the linear/translational velocity, and $\Omega$ is the angular velocity. While the results can be extended to any order, in this study, we only focus on the unicycle design to address objective functions that behave locally as a fourth-order degree polynomial in consistency with (grushkovskaya2025extremum, Assumption 1). Hence, we suppose:
@@
where $C_{1},C_{2}>0$. As analyzed, shown and simulated in grushkovskaya2025extremum, we need to excite the third-order Lie bracket for exponential convergence since the objective function order is of a fourth-degree. Hence, we choose our control law for $\mathrm{v}$ of the unicycle based on condition C3. That is,
@@
where $\varepsilon>0$ is a small parameter and $a,c>0$ are design constants. Now we are in a position to put all elements of the proposed design together, which is depicted in Figure 1. In state space representation, including an optional high-pass filter (HPF) into the system (ECC_2024; elgohary2025model), the proposed design becomes:
@@ 3.2 Stability Analysis
In this subsection, we provide the stability analysis for the proposed unicycle design (6). We will follow the traditional methodology in Lie bracket-based ESC literature (e.g., durr2013lie; grushkovskaya2018class; pokhrel2023higher) where the stability property of the ESC system is characterized by the corresponding Lie bracket system (LBS). The corresponding LBS to (6) without the optional HPF (i.e., $h=0$) is a third-order LBS based on condition C3 of the form (grushkovskaya2025extremum):
@@
For any initial condition $\bm{X}_{0}=\bm{X}(0)=\bm{\bar{X}}(0)$, from Theorem 1, there exists $q_{1}>0$ and $q_{2}>0$ such that $\bm{\bar{X}}-\bm{X}_{d}||\leq q_{1}e^{-q_{2}t}$. Moreover, per (pokhrel2023higher, Theorem 4), there exists $d(\varepsilon)\to 0$ as $\varepsilon\to 0$ such that $\bm{X}-\bm{\bar{X}}||\leq d(\varepsilon)$ for all $T>0$ and $t\in[0,T]$. Hence, the inequality (13) becomes:
@@
The inequality (14) guarantees that for any positive time $T$, which can be made large as needed, the trajectories of the unicycle system (6) will decay exponentially to a neighborhood about the extremum $\bm{X}_{d}$. Said neighborhood can be made arbitrarily small via the parameter $\varepsilon$. The reader can refer to khalil2002; Maggia2020higherOrderAvg; pokhrel2023higher for more details on the concept of practical exponential stability.
@@ 3.3 Tuning Control Parameters
For tuning the exponentially convergent unicycle system, we provide some basic guidelines for the three control parameters that need to be tuned: $c$, $a$, and $\varepsilon$. The parameter $c$ acts as a learning rate parameter that is scaling the feedback of the objective function measurement and is affecting the rate of the velocity $v$. In simulation, better flexibility can be expected for $c$ as one can theoretically increase, as needed, the rate by which the feedback measurement of $J$ affects the rate of $v$. However, in real-world applications, there is a hardware limitation for how $v$ can change (i.e., limitation for the rate of $v$), hence $c$ can be increased carefully. The parameter $a$ is a scale of the added perturbation signal that is needed for Lie bracket realization. In simulations, balance between $a$ and $c$ can be considered. However, in real-world experiments, $a$ can be chosen such that it does not cause hardware violation for the rate of $v$, so $a$ can always be chosen after figuring out the viable value(s) of $c$. As far as $\varepsilon$ is concerned, in theory/simulations the smaller the $\varepsilon$ the better; however, this is only possible in theory/simulations and is limited by the practical perturbation frequency the hardware implementation can admit. Furthermore, we note that the selection of $c$ and $a$ must be in accordance with Theorem 1, such that either condition $\bm{i}$ or $\bm{ii}$ is satisfied. Lastly, when selecting $\varepsilon$, it must be sufficiently small and less than the threshold, $\varepsilon^{*}$, where $\varepsilon^{*}$ is not explicitly known in averaging analysis context (Maggia2020higherOrderAvg). We note that the relationship between the perturbation frequency $\omega$ and $\varepsilon$ is $\omega=\frac{2\pi}{\varepsilon}$.
@@ 4.1 Experimental Setup
For experiments where the objective function is known, a motion capture system (MCS) is used to track the robot’s position in the $(x,y)$ coordinates. For the light source-seeking experiment, an analog light sensor connected to an Arduino Nano ESP32 board is mounted on the robot to measure light intensity. The robot used in our experiment is TurtleBot3 (see detailed information about this robot in bajpai2024investigating; ECC_2024). The sensor readings are transmitted to the computer and used as feedback for the ESC system. The MCS is also used in this case to record and observe the robot’s position for performance evaluation. The light sensor consists of a small photoresistor that measures the intensity of light incident on its surface. As the light intensity increases, the sensor output decreases; hence, minimizing the sensor’s measured value corresponds to approaching the point of maximum light intensity.
Figure 2: Modeling, Dynamics, and Control Lab (MDCL_Lab): (1) light source; (2) TurtleBot3 robot; (3) motion capture system (MCS); (4) analog light sensor; (5) Arduino Nano ESP32; (6) MCS markers.

### v3

@@ Abstract
This paper introduces a novel model-free, real-time unicycle-based source seeking design. This design autonomously steers the unicycle dynamic system towards the extremum point of an objective function or physical/scalar signal that is unknown expression-wise, but accessible via measurements. A key contribution of this paper is that the introduced design converges exponentially to the extremum point of objective functions (or scalar signals) that behave locally like a higher-degree power function (e.g., fourth-degree polynomial function) as opposed to locally quadratic objective functions, the usual case in literature. We provide theoretical results and design characterization, supported by a variety of simulation results that demonstrate the robustness of the proposed design, including cases with different initial conditions and measurement delays/noise. Also, for the first time in the literature, we provide experimental robotic results that demonstrate the effectiveness of the proposed design and its exponential convergence ability. These experimental results confirm that the proposed exponentially convergent extremum seeking design can be practically realized on a physical robotic platform under real-world sensing and actuation constraints.
@@ 1 Introduction
Extremum seeking control (ESC) techniques are model-free, real-time control methods that drive a system to the optimum of a given objective function (Ariyur and Krstic (2003); Tan et al. (2010); Guay and Dochain (2015); Scheinker (2024)), which may be unknown, provided it can be measured. ESC schemes are particularly attractive across various fields due to their minimal information requirements: they rely solely on applying perturbations to system parameters or inputs, and measuring the corresponding outputs of the objective function. Using these measurements in a feedback loop, ESC algorithms iteratively adjust the system input or parameter values to drive it towards the extremum, as can be seen in many applications, e.g., Krstic and Cochran (2008); Dochain et al. (2011); Calli et al. (2012); Bajpai et al. (2024); Elgohary et al. (2025); Elgohary and Eisa (2025a, b); Pokhrel and Eisa (2022); Eisa and Pokhrel (2023); Grushkovskaya et al. (2018a); Grushkovskaya and Zuyev (2024).
@@
One conventional problem that has increasingly involved the use of ESC methods is source-seeking (Bajpai (2024); Ghods (2011); Zhu et al. (2013); Li et al. (2014); Bulgur et al. (2018)). Source-seeking is the problem of steering a system (e.g., a mobile robot) autonomously to the maximum or minimum intensity of a physical/scalar signal present in a given domain using only on-board sensor measurements (e.g., heat, light, chemical concentration, among others). The advantage of ESC lies in the fact that it does not require knowledge of the signal’s distribution or explicit modeling of the system to drive the system, autonomously, to the extremum point. Due to the model-free nature of ESC methods, source-seeking techniques based on ESC utilized a simple, basic unicycle model to represent the system within the control design (Pokhrel et al. (2024); Matveev et al. (2011); Khong et al. (2014); Cochran et al. (2009); Grushkovskaya et al. (2018a); Suttner and Krstić (2022); Yilmaz et al. (2025); Todorovski and Krstić (2024); Elgohary et al. (2025)). As a result, unicycle-based methods for source-seeking often resort to Lie bracket averaging (Dürr et al. (2013); Scheinker (2017); Grushkovskaya et al. (2018b); Ghadiri-Modarres and Mojiri (2020); Pokhrel and Eisa (2023b)) in theoretical analysis and design. It is worth noting that the aforementioned unicycle-based designs for source-seeking are based on ESC laws that are gradient-based, which is more effective (e.g., guarantee exponential convergence) if the objective function or the scalar signal behaves locally like a quadratic function.
@@
On the other hand, optimization problems involving fourth-order polynomial functions have become an increasingly prevalent topic in the literature (Qi et al. (2004); Zhang and Wu (2015)). In fact, some emerging applications in science and engineering involve the need to optimize and control a fourth-order polynomial function, such as but not limited to the distribution of the coefficient of performance for power optimization ($C_{p}$) in wind turbines (Eisa (2019)) and pure-quartic solitons in optics, which exhibit fourth-order dispersion and have possible applications with lasers (Soltani et al. (2023)).
@@
Motivation and Contribution. In our recent work Grushkovskaya and Eisa (2025), new ESC laws were derived based on higher-order Lie bracket averaging. This paved the way for the prospect of having exponentially convergent ESC laws for objective functions that behave locally not as quadratic functions, but as power functions with higher-order degrees. Inspired by the prospect of realizing exponential convergence for high-order objective functions, we seek to utilize the ESC method in Grushkovskaya and Eisa (2025) and propose a novel, first-of-its-kind, exponentially convergent unicycle-based design for source-seeking. The main contribution of this work is the development and experimental validation of an exponentially convergent unicycle-based ESC law that realizes higher-order Lie bracket motion in real-time. In this preliminary study, we only focus on source-seeking of objective functions that behave locally as a fourth-degree power function. We prove stability of the proposed design and provide a variety of simulation results that demonstrate the robustness of the proposed design, including cases with different initial conditions and measurement delays/noise. Additionally, we provide experimental results using a TurtleBot3 robot (similar to Bajpai et al. (2024); Elgohary et al. (2025)) and compare the proposed third-order Lie bracket-based design with traditional first-order Lie bracket-based designs. We also perform a light source-seeking experiment in a complete model-free fashion. Our results illustrate the potential and effectiveness of the proposed exponentially convergent unicycle design for source-seeking.
@@ 2 Brief Background
In this section, we provide brief context regarding the exponentially convergent ESC law that we use for higher-order objective functions. We consider the two-input control-affine ESC of the following form, as introduced in Grushkovskaya and Eisa (2025):
@@
where $x\in D\subset\mathbb{R}^{n}$ denotes the system’s state, $J:D\to\mathbb{R}$ is the objective function, $u_{1}^{\varepsilon}(t)=\varepsilon^{1/N-1}v_{1}(t/\varepsilon)$, $u_{2}^{\varepsilon}(t)=\varepsilon^{1/N-1}v_{2}(t/\varepsilon)$, $N-1\in\mathbb{N}$ denotes the order of the Lie bracket to be excited, and $\varepsilon>0$ is a small positive parameter, such that the dithers $v_{1}(t/\varepsilon)$ and $v_{2}(t/\varepsilon)$ excite the Lie bracket
@@
Moreover, we adopt all mathematical assumptions on the vector fields $g_{1}$, $g_{2}$ provided in (Grushkovskaya and Eisa, 2025, Section II) and recall that $g_{1},g_{2}:\mathbb{R}\to\mathbb{R}^{n}$ are smooth vector fields evaluated at the scalar argument $J(x)$. However, we state below the assumption clarifying the structure and class of objective functions admitted in Grushkovskaya and Eisa (2025) as well as in this paper:
@@
Assumption A1.1 states that $J$ has an isolated minimum, $x^{*}$, at which it has a value $J^{*}$. Assumption A1.2 provides the requirement that the admitted objective function, $J$, behaves locally as a power function of degree $m$; this is consistent with ESC literature (e.g., Grushkovskaya et al. (2018b)).
The selection of the dither signals is done as in Grushkovskaya and Zuyev (2024); Grushkovskaya and Eisa (2025), conditions C1–C3:
@@
From Grushkovskaya and Eisa (2025), $g_{1},\ g_{2}$ are chosen according to:
@@
with the simple selection to satisfy (2) being $g_{1}(z)=-1^{(N+1)}z$ and $g_{2}(z)=1$. It is also to be noted that the conditions C1–C3 above guarantee that only the desired Lie bracket is excited while all others vanish per Gauthier and Kawski (2014); Pokhrel and Eisa (2023b); Grushkovskaya and Zuyev (2024). In this paper, we are interested in third-order Lie bracket excitation (i.e., condition C3).
@@ 3.1 Proposed Design
We aim to use an ESC law based on condition C3, provided in the previous section, to achieve an exponentially convergent unicycle design. Let us consider the kinematic differential equations for unicycle dynamics with constant angular velocity as follows (Elgohary et al. (2025); Grushkovskaya et al. (2018a)):
@@
where $x$ and $y$ describe the current position/coordinates, $\mathrm{v}$ is the linear/translational velocity, and $\Omega$ is the angular velocity. While the results can be extended to any order, in this study, we only focus on the unicycle design to address objective functions that behave locally as a fourth-order degree polynomial in consistency with (Grushkovskaya and Eisa, 2025, Assumption 1). Hence, we suppose:
@@
where $C_{1},C_{2}>0$. As analyzed, shown and simulated in Grushkovskaya and Eisa (2025), we need to excite the third-order Lie bracket for exponential convergence since the objective function order is of a fourth-degree. Hence, we choose our control law for $\mathrm{v}$ of the unicycle based on condition C3. That is,
@@
where $\varepsilon>0$ is a small parameter and $a,c>0$ are design constants. Now we are in a position to put all elements of the proposed design together, which is depicted in Figure 1. In state space representation, including an optional high-pass filter (HPF) into the system (Bajpai et al. (2024); Elgohary et al. (2025)), the proposed design becomes:
@@ 3.2 Stability Analysis
In this subsection, we provide the stability analysis for the proposed unicycle design (6). We will follow the traditional methodology in Lie bracket-based ESC literature (e.g., Dürr et al. (2013); Grushkovskaya et al. (2018b); Pokhrel and Eisa (2023b)) where the stability property of the ESC system is characterized by the corresponding Lie bracket system (LBS). The corresponding LBS to (6) without the optional HPF (i.e., $h=0$) is a third-order LBS based on condition C3 of the form (Grushkovskaya and Eisa (2025)):
@@
For any initial condition $\bm{X}_{0}=\bm{X}(0)=\bm{\bar{X}}(0)$, from Theorem 1, there exists $q_{1}>0$ and $q_{2}>0$ such that $\bm{\bar{X}}-\bm{X}_{d}||\leq q_{1}e^{-q_{2}t}$. Moreover, per (Pokhrel and Eisa, 2023b, Theorem 4), there exists $d(\varepsilon)\to 0$ as $\varepsilon\to 0$ such that $\bm{X}-\bm{\bar{X}}||\leq d(\varepsilon)$ for all $T>0$ and $t\in[0,T]$. Hence, the inequality (13) becomes:
@@
The inequality (14) guarantees that for any positive time $T$, which can be made large as needed, the trajectories of the unicycle system (6) will decay exponentially to a neighborhood about the extremum $\bm{X}_{d}$. Said neighborhood can be made arbitrarily small via the parameter $\varepsilon$. The reader can refer to Khalil and Grizzle (2002); Maggia et al. (2020); Pokhrel and Eisa (2023b) for more details on the concept of practical exponential stability.
@@ 3.3 Tuning Control Parameters
For tuning the exponentially convergent unicycle system, we provide some basic guidelines for the three control parameters that need to be tuned: $c$, $a$, and $\varepsilon$. The parameter $c$ acts as a learning rate parameter that is scaling the feedback of the objective function measurement and is affecting the rate of the velocity $v$. In simulation, better flexibility can be expected for $c$ as one can theoretically increase, as needed, the rate by which the feedback measurement of $J$ affects the rate of $v$. However, in real-world applications, there is a hardware limitation for how $v$ can change (i.e., limitation for the rate of $v$), hence $c$ can be increased carefully. The parameter $a$ is a scale of the added perturbation signal that is needed for Lie bracket realization. In simulations, balance between $a$ and $c$ can be considered. However, in real-world experiments, $a$ can be chosen such that it does not cause hardware violation for the rate of $v$, so $a$ can always be chosen after figuring out the viable value(s) of $c$.
As far as $\varepsilon$ is concerned, in theory/simulations, the smaller the $\varepsilon$, the better; however, this is only possible in theory/simulations and is limited by the practical perturbation frequency the hardware implementation can admit. In particular, $\varepsilon$ determines the perturbation frequency $\omega=\frac{2\pi}{\varepsilon}$, and excessively small values of $\varepsilon$ correspond to high-frequency dither signals that may exceed actuator bandwidth, violate the Nyquist limit associated with the fixed control-loop sampling rate, or induce additional computational delays. Consequently, $\varepsilon$ is selected as the smallest value that can be reliably implemented while maintaining stable command tracking.
Furthermore, we note that the selection of $c$ and $a$ must be in accordance with Theorem 1, such that either condition $\bm{i}$ or $\bm{ii}$ is satisfied. Lastly, when selecting $\varepsilon$, it must be sufficiently small and less than the threshold, $\varepsilon^{*}$, where $\varepsilon^{*}$ is not explicitly known in averaging analysis context (Maggia et al. (2020)). Additionally, in practice $\varepsilon$ may be selected by first identifying an admissible perturbation frequency $\omega$ based on hardware and sampling constraints, and then computing $\varepsilon=\frac{2\pi}{\omega}$.
@@ 4.1 Experimental Setup
For experiments where the objective function is known, a motion capture system (MCS) is used to track the robot’s position in the $(x,y)$ coordinates. For the light source-seeking experiment, an analog light sensor connected to an Arduino Nano ESP32 board is mounted on the robot to measure light intensity. The robot used in our experiment is TurtleBot3 (see detailed information about this robot in Bajpai (2024); Bajpai et al. (2024)). The sensor readings are transmitted to the computer and used as feedback for the ESC system. The MCS is also used in this case to record and observe the robot’s position for performance evaluation. The light sensor consists of a small photoresistor that measures the intensity of light incident on its surface. As the light intensity increases, the sensor output decreases; hence, minimizing the sensor’s measured value corresponds to approaching the point of maximum light intensity.
Figure 2: Modeling, Dynamics, and Control Lab (MDCL (2025)): (1) light source; (2) TurtleBot3 robot; (3) motion capture system (MCS); (4) analog light sensor; (5) Arduino Nano ESP32; (6) MCS markers.